OpenStudy (anonymous):

What is the fourth term of (d + 4b)3?

jhonyy9 (jhonyy9):

3 is exponent ?

OpenStudy (anonymous):

Yes,i meant to put ^3 sorry

OpenStudy (cwrw238):

(r + 1)th term of (a + x)^n = nCr a^(n-r) x^r

OpenStudy (cwrw238):

for 4th term of your expression r = 3, a = d, x = 4b and n = 3

OpenStudy (turingtest):

to state @cwrw238 's answer another way: \[(a+b)^n=\sum_{k=0}^n\binom nka^{n-k}b^k\]where\[\binom nk=\frac{n!}{k!(n-k)!}\]for the \(4^{th}\) term that means that \(k=3\)

OpenStudy (anonymous):

are you satisfied with the explanations they gave you or do you want an alternative explanation using another method?